# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v3

@@ 2.3. Overview of Sisyphus
To address these limitations, we propose Sisyphus, which focuses on linear code to offer a space capable of exploring only legal code transformations, pragma insertions, and tile size selections as a single optimization problem. Leveraging a Non-Linear Programming (NLP) approach, Sisyphus can efficiently explore this space, benefiting from accurate modeling enabled by compile-time analysis feasible for linear code. The space comprises three levels of tiling, each corresponding to a specific optimization: fine-grained unrolling, pipelining, and coarse-grained unrolling, with loops executed sequentially and data transferred between off-chip and on-chip.
For gemm, Sisyphus excels in boosting parallelism by splitting and permuting loops, all without depending on the reduction loop. In this particular setup, the NLP discovers the configuration showcased in Listing 3.
1// possibilities to cache on-chip A, B and C
2/********* Level 0 *********/
3for (int i0 = 0; i0 < 1; i0++)
4// possibilities to cache on-chip C
5 for (int j0 = 0; j0 < 1; j0++)
6// possibilities to cache on-chip C
7/********* Level 1 *********/
8 for (int i1 = 0; i1 < 1; i1++)
9 for (int j1 = 0; j1 < 55; j1++)
10#pragma HLS pipeline
11/********* Level 2 *********/
12 for (int i2 = 0; i2 < 200; i2++)
13#pragma HLS unroll
14 for (int j2 = 0; j2 < 4; j2++)
15#pragma HLS unroll
16 C[i0*200+i2][j0*220+j1*4+j2] *=beta ;
17/********* Level 0 *********/
18for (int i0 = 0; i0 < 1; i0++)
19// possibilities to cache on-chip A and C
20 for (int j0 = 0; j0 < 1; j0++)
21// possibilities to cache on-chip B and C
22 for (int k0 = 0; k0 < 60; k0++)
23// possibilities to cache on-chip A and B
24/********* Level 1 *********/
25 for (int i1 = 0; i1 < 1; i1++)
26 for (int k1 = 0; k1 < 1; k1++)
27 for (int j1 = 0; j1 < 220; j1++)
28#pragma HLS pipeline
29/********* Level 2 *********/
30 for (int i2 = 0; i2 < 200; i2++)
31#pragma HLS unroll
32 for (int j2 = 0; j2 < 1; j2++)
33#pragma HLS unroll
34 for (int k2 = 0; k2 < 4; k2++)
@@ 16 C[i0*200+i2][j0*220+j1*4+j2] *=beta ;
36 i = i0*200+i2; j = j0*220+j1*1+j2;
37 k = k0 * 4 + k2;
38 C[i][j] +=alpha * A[i][k] * B[k][j];
Listing 3: Code of gemm found by Sisyphus
Both pipelined loops feature an II of 1, enabling a throughput of 210 GF/s. Partially unrolling the reduction loop allows for parallel execution of the two multiplications.
Concerning CNN, the NLP swiftly identifies a solution within 4.5 minutes. This solution optimizes tile configuration, caches a data tile of the array on-chip, partially unrolls one reduction loop and two non-reduction loops, and pipelines a non-reduction loop. As a result, the design achieves a throughput of 339 GF/s. The design discovered by Sisyphus is showcased in Listing 4.
1for (int i0 = 0; i0 < 16; i0++) {
2 load_weight(weight,vweight,i0); // size 16*256*5*5
3 load_output(output, voutput, i0); // size 16*224*224
4 for (int j0 = 0; j0 < 256; j0++) {
5 load_input(input, vinput, j0); // size 228*228
6 for (int q0 = 0; q0 < 5; q0++)
7 for (int w0 = 0; w0 < 14; w0++)
8 for (int h1 = 0; h1 < 224; h1++)
9#pragma HLS pipeline
10 for (int i2 = 0; i2 < 16; i2++) // unrolled
11 for (int w2 = 0; w2 < 16; w2++) // unrolled
12 for (int p2 = 0; p2 < 5; p2++){ // unrolled
13 i = ...
14 output[i2][h][w] +=weight[i2][j][p][q]
15 * input[j2][h+p][w+q];}}
16 store_output(voutput, output, i0);}
Listing 4: Code of CNN found by Sisyphus
3. Unified Space
We proceed to develop a methodology that integrates code transformation, tile size selection and pragma insertion into a unified optimization framework. This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously. We illustrate our code transformation process step by step using the example provided in Listing 1.
Maximum Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement. This distribution strategy creates ample opportunities for parallelization by maximizing the separation between dependencies. Typically, this results in one statement per loop body with perfectly nested loops. To achieve this, we employ ISCC (Verdoolaege, 2011). It allows us to explore various schedule distributions and validate the legality of transformations by ensuring dependency constraints are preserved. After distribution, the example (Listing 1) yields two loop bodies. The first statement iterates through loops $i$ and $j$, while the second statement iterates through loops $i$, $k$, and $j$.
Strip-mining  Following maximal distribution, we perform strip-mining on each loop twice. This process replaces the original loop, with a trip count $TC\_I0$, with three new loops, each having trip counts $TC\_I0\_0,TC\_I0\_1$ and $TC\_I0\_2$ with $TC\_I0\_0\times TC\_I0\_1\times TC\_I0\_2=TC\_I0$. Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation. Subsequently, we can arrange the loops in different permutations if it is legal. Following the implementation of three levels of strip-mining, the first loop body of the code from Listing 1 undergoes transformation into the code depicted in Listing 5.
1for (i0 = 0; i0 < TC_I0_0; i0++)
2 for (i1 = 0; i1 < TC_I0_1; i1++)
3 for (i2 = 0; i2 < TC_I0_2; i2++)
4 for (j0 = 0; j0 < TC_J0_0; j0++)
5 for (j1 = 0; j1 < TC_J0_1; j1++)
6 for (j2 = 0; j2 < TC_J0_2; j2++)
7 i = i0 * TC_I0_1 * TC_I0_2 + i1 * TC_I0_2 + i2;
8 ...
9 C[i][j] *= beta;
Listing 5: Code of the first loop body of gemm with three level strip mining
Following the explanation provided in the next paragraph, strip mining allows us to create opportunities for fine-grained unrolling, pipelining, sequential execution, coarse-grained unrolling and tiling if loop permutation is legal. This expands the range of possibilities, allowing each loop to integrate diverse hardware directives and be reorganized across different levels of the code.
Loop Permutation  Next, we consider possible permutations of the strip-mined loops. If these permutations are legal, we establish three loop levels. We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011).
The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize tree reduction if a reduction loop is unrolled and the option is enabled. The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations. Finally, the outermost level coordinates coarse-grained unrolling, tiling, and/or sequential execution, optimizing to increase the parallelism, and controlling the size of the on-chip buffer through tiling. If the loop body consists of only one statement and loop permutation is feasible, we disable coarse-grained unrolling. This will achieve equivalent results because our template permits unrolling the same loop innermost for fine-grained unrolling. Furthermore, coarse-grained unrolling entails duplicating each array, including read-only arrays, to ensure full overlapping of the modules, even when there are no dependencies between the concurrently executed modules.
In scenarios where permutations enable us to achieve perfectly nested loops capable of being pipelined, only one loop can usually be pipelined. Trying to pipeline all loops simultaneously, akin to pragma loop_pipelined, significantly complicates the design and extends synthesis times unreasonably. Thus, if a loop in the middle level is pipelined, it implies that the trip counts of the other loops (in the middle level) are reduced to one, ensuring manageable design complexity and synthesis times. As explained earlier, loop order selection is not necessary at the pipeline level. In the fine-grained level, arranging loops is straightforward; placing the reduction loop innermost enables efficient tree reduction. Conversely, at the coarse-grained or tiled level, prioritizing memory transfer latency is key. Therefore, we opt for a loop order that maintains the output stationary, minimizing costs associated with loading on-chip and storing off-chip during tile iterations, especially when the output is not fully transferred to the chip.
In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations. It is notable that our framework maintains the original loop order. In Listing 3, if the loop on line 12 ($i2$) iterates once and the loop on line 3 ($i0$) iterates 200 times, we keep the original loop order, then proceed to pipeline and partially unroll loop $j$.
Fusion  We have opted not to incorporate fusion into our model for several reasons. Firstly, our objective is to maximize parallelization for computation-bound kernels. Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities. Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation. Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times.
Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $TC\_I0\_0$). Additionally, the design must adhere to resource constraints, including the usage of DSP and on-chip memory. Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources. The on-chip buffers must not exceed the on-chip memory capacity. Consequently, for a large problem size, the outermost level will enable control over the buffer sizes. Thus, the space we consider encompasses the original schedule on which we can insert pragmas, but it also includes all the transformations mentioned previously for which the pragmas are inserted.
@@ 4. NLP
Presented here is a comprehensive set of constraints and variables employed in a non-linear program aimed at discovering the theoretical solution space outlined in Section 3. We employ the methodology proposed by (Pouget et al., 2024b, a), adapted to our specific context. Similar to their approach, we establish an estimation of the latency by assuming optimistic DSP utilization, assuming perfect resource reuse. However, we afford the user the flexibility to adjust the DSP limit or opt for a pessimistic DSP utilization scenario, where no reuse between statements is considered. Moreover, users have the ability to adjust the size of on-chip memory, define the maximum number of array partitions, and customize the latency and resources allocated for each operation. This flexibility ensures adaptability across various platforms and compilers.
To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
@@ 4.1. Variables
Tables 1 and 2 delineate the sets, variables, and constants utilized in our NLP formulation.
@@
$\mathcal{L}_{s}^{red}$ the set of reduction loops which iterate the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops which iterates the array $a$ at the dimension $d$
@@
$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$
@@
Table 2. Overview of the variable employed in formulating the NLP
4.2. Constraints
Now, we explore the precise meaning and implications of each constraint.
Trip Count  Equation 1 constrain the trip count of each loop, ensuring that the product of the trip counts equals the original trip count.
(1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
Coarse-grained unrolling  We enforce coarse-grained unrolling solely for non-reduction loops (Eq. 4), where the unroll factor (UF) must divide and be less than or equal to the trip count of the current loop (Eq. 5 and 2). As elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq. 3).
@@ 4.2. Constraints
Pipeline  Constraints 6, 7, 8, and 9 facilitate the selection of loop pipelining, allowing only one pipeline per nested loop and computing the initiation interval (II) in accordance with the chosen pipelined loop. The II is determined based on the loop’s properties and the iteration latency of the reduction operation, specifically when the loop being pipelined is a reduction loop, following the approach outlined in (Pouget et al., 2024b, a).
(6) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
(7) $\forall l\in\mathcal{L},(1-loop_{l}\_pip)\times TC_{l,1}==1$
On-chip memory  Constraints 10 and 11 ensure that each array is cached on-chip at a single location within the code, and that the on-chip tiled data fits within the available on-chip memory.
@@
Array partitioning  Equations 12, 13 and 14 guarantee that the maximum array partitioning is not exceeded, and it ensures that the partitioning of the array is greater than or equal to the maximum unroll factor applied to that dimension. Thanks to Equation 13, which ensures that each unroll factor ($TC_{l,2}$) divides $AP_{a,d}$, $AP_{a,d}$ directly determines the array partitioning applied in the generated code.
@@
DSP utilization  Lastly, Equations 15, 16, 17, and 18 compute the DSP utilization under optimistic (full reuse) and pessimistic (no reuse) DSP reuse assumptions.

### v4

@@ 2.3. Overview of Sisyphus
To overcome these limitations, we introduce Sisyphus, a solution tailored for affine programs, where loop bounds and conditionals are affine functions of surrounding loop iterators and parameters (commonly referred to as Static Control Parts in the Polyhedral Model) (Girbal et al., 2006; Feautrier, 1988). Our objective is to define a search space that allows for the selection of various parameters, such as pragma insertion, loop ordering and tiling, within a single optimization problem. This approach avoids the need for backtracking or handling disjoint optimization problems. Our framework efficiently and rapidly explores this large design space, overcoming the limitations of other frameworks that either restrict the search space too much or explore a broad space with a DSE that only partially covers it. Most importantly, it guarantees that only valid code transformations are considered. To accomplish this, we propose a two-step process: first, the creation of a space tailored to each kernel, followed by the exploration of that space. The first step involves creating a template with fully distributed code that includes three levels of loops, each targeting specific optimizations. To achieve this, we strip-mined each original loop twice, ensuring that one of these loops is included at each level (assuming the permutation is valid). As a result, each original loop can be utilized across all three levels. For instance, in Listing 1, the loop $k$ (L2) is used in level 0 alongside loop L5 and in level 2 with loop L11 in Listing LABEL:lst:small_mem. Specifically, we have an innermost level for fine-grained unrolling, a level dedicated to pipelining, and another level focused on on-chip data caching based on available resources. As demonstrated in Listings LABEL:lst:small_mem and 2, depending on the resource constraints specified by the user, we obtain two different code versions. For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (line 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints. Once this template is generated, it establishes a search space where the unknowns consist of the loop bounds, loop ordering, buffer sizes for on-chip arrays, and the locations in the code where data transfers take place. Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem. This approach allows us to efficiently navigate the theoretical space within seconds or minutes, leveraging accurate modeling made possible by compile-time analysis for affine programs. It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS, as done in AutoDSE, would be impractical due to the sheer size of the search space. In contrast, the NLP approach enables us to explore this space quickly and efficiently, allowing for a more comprehensive investigation than traditional methods. For gemm, Sisyphus significantly enhances parallelism by splitting and permuting loops without relying on the reduction loop. In this configuration, the NLP identifies the setup shown in Listing 2. Both pipelined loops achieve an initiation interval (II) of 1, resulting in a throughput of 210 GFLOPs per seconds (GF/s). The reduction loop (L11) is partially unrolled, allowing the multiplications to be executed in parallel while performing four additions sequentially. This small unrolling, selected by the NLP, deepens the pipeline, improving the quality of results (QoR) even though the reduction must remain sequential. For the CNN, the NLP selects all parameters within the design space, including loop order after exploring 6! possibilities, data transfer points, on-chip array size, and the level of parallelism. This comprehensive analysis takes just 13 minutes and yields a throughput of 341 GF/s. Thanks to a well-defined space that only considers legal transformations and exploration guided by a cost model formulated as an NLP, our framework is able to select the theoretical optimum in the entire space in a matter of seconds or minutes and generate a design.
3. Unified Space
We proceed to develop a methodology that integrates code transformation, tile size selection and pragma insertion into a unified optimization framework. This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously. We illustrate our code transformation process step by step using the example from Listings 1 to 2. Input  Sisyphus takes an affine C/C++ file as input and automatically extracts a polyhedral representation of the code using PoCC (PoCC, [n. d.]). This representation allows us to gather all relevant information from the code, such as loop order, loop bounds, array sizes, and more. We will use this information to create the template and to automatically generate the NLP discussed in Section 4. Our framework will be particularly effective for codes that are fully distributable and fully permutable. However, in cases where the transformations described later are not legally possible, our framework still functions but restricts the solution space more. Maximal Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement. This distribution strategy creates ample opportunities for parallelization by maximizing the distance between dependencies (reverse of fusion). Typically, this results in one statement per loop body with perfectly nested loops. To achieve this, we employ ISCC (Verdoolaege, 2011). It allows us to explore various schedule distributions and validate the legality of transformations by ensuring dependency constraints are preserved. The code presented in Listing LABEL:lst:distributed displays the fully distributed version of the code from Listing 1. Strip-mining  After maximal distribution, we apply strip-mining to each loop twice. This process transforms the original loop, such as the loop L0 in Listing LABEL:lst:distributed, into three loops with trip counts $I0\_S0$, $I1\_S0$, and $I2\_S0$ (L0, L1, L2 in Listing LABEL:lst:strip). The product of these three loop trip counts equals the original trip count, i.e., $I0\_S0\times I1\_S0\times I2\_S0=200$. Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation. Subsequently, we can arrange the loops in different permutations if it is legal. After applying three levels of strip-mining, the code from Listing LABEL:lst:distributed is transformed into the code seen in Listing LABEL:lst:strip. Creation of different levels  Next, we consider possible permutations of the strip-mined loops. If these permutations are legal, we establish three loop levels. We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011). Level 2:  The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize a tree reduction if a reduction loop is unrolled and the option is enabled. For example, this is illustrated by loops L9, L10, and L11 in Listing 2. Level 1:  The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations. In cases where loop permutations allow us to achieve perfectly nested loops suitable for pipelining, we restrict pipelining to a single loop to ensure efficient synthesis. Flattening the loops at Level 1 into a single loop and then pipelining the resulting loop introduces significant design complexity and leads to excessively long synthesis times. Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one, ensuring manageable design complexity and synthesis times. For instance, in Listing 2, the loop $j$ (L8) is pipelined for the second statement (S1). This results in the loops $k$ and $i$ (L10 and L11 in Listing 3 ) having a trip count of 1, enabling them to be eliminated from the code. Level 0:  Finally, the outermost level manages coarse-grained unrolling, tiling, and/or sequential execution. It not only controls parallelism at a coarse level but also determines the on-chip buffer size through tiling. As described in the following paragraph, the loop order at this level is included in the design space. This level is illustrated by loops L5, L6, and L7 in Listing LABEL:lst:small_mem. The loop bounds determine the on-chip buffer sizes for arrays A and B, transferred below loop L5. Resource constraints also influence the sequential execution of loop L7. If the loop body contains a single statement and loop permutation is feasible, we disable coarse-grained unrolling. This achieves equivalent results since our template allows fine-grained unrolling of the innermost loop. Loop Permutation  Loop order selection is not required for levels 1 and 2, as only a single loop is pipelined, and the innermost level (level 2) is fully unrolled, making the loop order irrelevant. Conversely, at the outermost level (level 0), all permutations are considered, and the NLP selects the permutation. In Listings LABEL:lst:small_mem and 2, we can see that the NLP selected two different permutations for the statement S1 (loops L5, L6, and L7).
1int perm_S0[2] // Permutation order (0 for i0, 1 for j0)
2int perm_S1[3] //Array to store permutation order (i0,j0,k0)
3int lb_S0[2] = {I0_S0, J0_S0}; // loop bounds for S0
4int lb_S1[3] = {I0_S1, J0_S1, K0_S1}; // loop bounds for S1
5// possibilities to cache on-chip A, B and C
6/********* Level 0 of S0 *********/
7for (L0_S0=0; L0_S0 < lb_S0[perm_S0[0]]; L0_S0++) // L0
8// possibilities to cache on-chip C
9 for (L1_S0=0; L1_S0 < lb_S0[perm_S0[1]]; L1_S0++) // L1
10 // possibilities to cache on-chip C
11/********* Level 1 of S0 *********/
12 for (j1=0; j1 < J1_S0; j1++) // L2
13 for (i1=0; i1 < I1_S0; i1++) // L3
14#pragma HLS pipeline
15/********* Level 2 of S0 *********/
16 for (i2=0; i2 < I2_S0; i2++) // L4
17#pragma HLS unroll
18 for (j2=0; j2 < J2_S0; j2++) // L5
19#pragma HLS unroll
20 C[i][j]*=beta; // S0
21/********* Level 0 of S1 *********/
22for (L0_S1=0; L0_S1 < lb_S1[perm_S1[0]]; L0_S1++) // L6
23// possibilities to cache on-chip A, B and C
24 for (L1_S1=0; L1_S1 < lb_S1[perm_S1[1]]; L1_S1++) // L7
25 // possibilities to cache on-chip A, B and C
26 for (L2_S1=0; L2_S1 < lb_S1[perm_S1[2]]; L2_S1++) // L8
27 // possibilities to cache on-chip A, B and C
28/********* Level 1 of S1 *********/
29 for (j1=0; j1 < J1_S1; j1++) // L9
30 for (k1=0; k1 < K1_S1; k1++) // L10
31 for (i1=0; i1 < I1_S1; i1++) // L11
32#pragma HLS pipeline
33/********* Level 2 of S1 *********/
34 for (i2=0; i2 < I2_S1; i2++) // L12
@@ 20 C[i][j]*=beta; // S0
36 for (j2=0; j2 < J2_S1; j2++) // L13
37#pragma HLS unroll
38 for (k2=0; k2 < K2_S1; k2++) // L14
39#pragma HLS unroll
40 i = i0 * I1_S1 * I2_S1 + i1 * I2_S1 + i2;
41 j = j0 * J1_S1 * J2_S1 + j1 * J2_S1 + j2;
42 k = k0 * K1_S1 * K2_S1 + k1 * K2_S1 + k2;
[…]
Listing 3: Template
Fusion  We have opted not to incorporate fusion into our model for several reasons. Firstly, our objective is to maximize parallelization for computation-bound kernels. Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities. Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation. Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times. Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations. The iterator of the loop retains its original name, with an added number indicating the level. For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining. Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized. Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $J0\_S0$) and the loop order of the level 0. Additionally, the design must adhere to resource constraints, including the usage of DSP and on-chip memory. Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources. The on-chip buffers size must not exceed the on-chip memory capacity. Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
@@ 4. NLP
Presented here is a comprehensive set of constraints and variables employed in a nonlinear program aimed at discovering the theoretical solution space outlined in Section 3. We employ the methodology proposed by (Pouget et al., 2024b, a), adapted to our specific context. Similar to their approach, we establish an estimation of the latency by assuming optimistic DSP utilization, assuming perfect resource reuse. However, we afford the user the flexibility to adjust the DSP limit or opt for a pessimistic DSP utilization scenario, where no reuse between statements is considered. Moreover, users have the ability to adjust the size of on-chip memory, define the maximum number of array partitions, and customize the latency and resources allocated for each operation. This flexibility ensures adaptability across various platforms and compilers. To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
@@ 4.1. Variables
Table 2 delineate the sets, variables, and constants utilized in our NLP formulation.
4.2. Constraints
Now, we explore the precise meaning and implications of each constraint. To make this more understandable, we provide examples using the listings presented in the paper. Trip Count  Equation 1 constrains the trip count of each loop, ensuring that the product of the trip counts equals the original trip count. The three loops after strip-mining, e.g., L6, L7, and L8 in Listing LABEL:lst:strip, must be equivalent to the original loop L2 in our example from Listing LABEL:lst:distributed. Therefore, we require $I0\_S0\times I1\_S0\times I2\_S0=200$.
(1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
Pipeline  Constraints 2, 3, 4, and 5 facilitate the selection of loop pipelining. Constraint 2 ensures that each loop has a boolean flag indicating whether the original loop is pipelined. For example, in Listings LABEL:lst:small_mem and 2, the statement $S1$ shows that the loop $j1$ (L8) is pipelined, while the original loop is $j$ (L3 in Listing 1), indicating that $loop_{j}\_pip=1$. As explained in Section 3, only one loop can be pipelined at each middle level. This means the sum of the boolean flags must be less than or equal to 1 (Eq. 4), and the trip count of all other non-pipelined loops at the middle level must be 1 (Eq. 3). For example, in the template (Listing 3), only loop L9 is pipelined at level 1 of S1 (loops L9, L10, and L11), so $loop_{j}\_pip=1$. Consequently, $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$ (Eq. 4), and Eq. 3 indicates that the trip counts of loops L10 and L11 are 1, allowing their removal.
@@ 4.2. Constraints
$\mathcal{L}_{s}^{red}$ the set of reduction loops iterating the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops iterating the array $a$ at dimension $d$
@@
$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$ (This number accounts for the number of times the statement is replicated due to unrolling)
@@
$perm_{s}$ ID of the permutation of the statement $s$
$burst_{a}$ The maximum bitwidth at which the array $a$ can be transferred from off-chip to on-chip
Table 2. Overview of the set, constant and variables employed in formulating the NLP
Eq. 5 allows the computation of the initiation interval (II), which is determined based on the loop’s properties and the iteration latency of the reduction operation. Specifically, when the loop being pipelined is a reduction loop, this follows the approach outlined in (Pouget et al., 2024b, a). In Listings LABEL:lst:small_mem and 2, the pipelined loops are non-reduction loops with no dependencies when iterating, hence $II=1$. However, if loop L10 in Listing 3 were pipelined, it would involve a reduction, meaning that at each iteration, C[i][j] needs to be read and written. Thus, to read for iteration $k+1$, we must wait for the write to finish at iteration $k$, which takes $IL_{+}$ cycles (latency of the reduction operation).
(2) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
(3) $\forall l\in\mathcal{L},(1-loop_{l}\_pip)\times TC_{l,1}==1$
Coarse-grained unrolling  We enforce coarse-grained unrolling solely for non-reduction loops (Eq. 8). For example, the loop L3 in Listing LABEL:lst:distributed cannot be coarse-grained unrolled as it is a reduction loop. Additionally, the unroll factor (UF) must divide the trip count of the current loop and be less than or equal to it (Eq. 9 and 6). However, as elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq. 7). This implies that for Listing 1 coarse-grained unrolling is disabled.
@@
On-chip memory  Constraints 10 and 11 ensure that each array, represented by a boolean (Eq. 10), is cached on-chip at a single location within the same loop body, as only one boolean can be true for each loop body (Eq. 11). In Listing 3, for instance, array $A$ can be cached either before any loops (line 5) or after L6, L7, or L8 (lines 23, 25, and 27). In Listing 2, $A$ is cached on-chip with the on-chip footprint matching the size of the original array (line 2). Meanwhile, in Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18, but with a smaller on-chip footprint compared to the original. Based on where each array is transferred, we can compute the on-chip footprint. For example, in Listing LABEL:lst:small_mem, array $A$ is transferred below loop L5, which has a trip count of 60, so the footprint of $A$ on-chip is $200\times\frac{240}{60}$.
@@
(11) $\forall a\in\mathcal{A},\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}},cache_{l}\_arr_{a}=1$
@@
Array partitioning  We consider unrolling only if all the data needed for unrolling can be accessed in parallel, i.e., these data are on different on-chip banks thanks to the pragma array_partitioning. For each array $a$ and each dimension $d$ of this array, we want the factor of the cyclic array partitioning for this dimension to be greater than the unroll factor (Eq. 15) to ensure that all data are on different banks. If the array is reused, we need to ensure that the array partitioning is optimal for each statement. Therefore, we force the array partitioning to be a multiple of each unroll factor that iterates through this dimension (Eq. 14). $AP_{a,d}$ gives the cyclic array partitioning factor for array $a$ in dimension $d$ that we use in the generated code. Eq. 13 assures that the array partitioning per array (product of the partitioning factors per dimension) is less than or equal to the limit fixed by the user. For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second dimension (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second loop body. Thus, we have $AP_{C,1}=200$ and $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, where $max_{part}$ is 1,024 in our case.
@@
DSP utilization  Finally, we need to manage DSP utilization. Given the challenges in predicting DSP reuse, we use a boolean flag, $reuse_{opt}$, set by the user to determine the level of DSP reuse. If $reuse_{opt}=1$, we assume an optimistic reuse scenario: DSPs used in one loop body can be directly reused by subsequent loop bodies if the operations are identical. However, if the operations differ, as discussed in (Pouget et al., 2024b, a), no reuse is assumed. Therefore, the total DSP usage is calculated as the sum of the maximum DSPs required for each operation across all loop bodies (Eq. 16). If $reuse_{opt}=0$, we assume no reuse between different loop bodies, and we sum the DSP requirements for each statement (Eq. 17). Based on the value of $reuse_{opt}$, the total DSP usage must be less than or equal to the available DSPs (Eq. 18 and 19). For instance, if a multiplication operation (*) uses 3 DSPs and an addition operation ($+$) uses 2 DSPs, in Listing 2, S0 requires $200\times 4\times 3$ DSPs for multiplications, and S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition. With optimistic reuse for the same operation (e.g., multiplication), the DSP usage is calculated as: $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$. With pessimistic reuse, the DSP usage is: $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.